### v5

# On the “verification” of a Nicolas inequality

Abstract.
Nicolas inequality we deal can be written as
(1) $e^{\gamma}\log\log N_{x}<\dfrac{N_{x}}{\varphi(N_{x})}\,$,
where $x\geq 2$, $N_{x}$ denotes the product of the primes less or equal than $x$, $\gamma$ is the Euler constant and $\varphi$ is the Euler totient function. We see that verification of (1) depends on the sign of the big-O function in the Mertens estimate for the sum of reciprocals of primes. Then, we analyze the sign of such an error term.
Key words and phrases:
Nicolas inequality, Mertens Theorem, big-O function, totient function
Note to the reader
First version of the present preprint has an error that brings us to an incorrect conclusion. In second version we amend this error, but we do not provide new contributions. In this new version we have corrected Step 1 of final discussion and added Remark 2.4, and also (as previously) we would advise the reader that the paper contains some imprecisions, that the author have left to the reader, and also could include some undetected mistakes, so the author appreciates any comment without attached files (in order to avoid fishing) to galdames@uv.es .
1. Introduction and notation
Nicolas inequality was introduced in [6, Theorem 2] and establish that one and only one of the following holds:
(T) Inequality (1) holds for every integer $x\geq 2$.
(F) Inequality (1) does not hold for an infinite set of integers $x$’s.
However, as far as we know, it is unsolved which of them is true.
Mertens proved in [5] that for $x\geq 2$
$\sum_{p\leq x}\frac{1}{p}=\log\log x+b+O(1/\log x)\,$,
where $b$ is sometimes called Mertens constant.
It turns out that if we apply adequately this Mertens estimate to the Nicola’s inequality we see that the sign of the error term $O(1/\log x)$ is decisive for the compliance or the failure of Nicolas inequality. In fact, we prove that if $O(1/\log x)$ is bounded from below by a positive number, then (T) holds, unlike if $O(1/\log x)$ is bounded from above by a negative number, then (F) holds. After that, follow the proof of Mertens Theorem and find that we cannot determine the sign of $O(1/\log x)$ with the known estimates. For the sake of simplicity we use the notation $N_{x}=\prod_{p\leq x}p$, instead of the original $N_{k}$, which is the product of the first $k$ primes. Notice that both are equivalent, because $N_{x}=N_{[x]}$, where $[x]$ is the floor function, and for any positive integer $k$ there is a real number $x$ such that $k=[x]$.
Let us introduce some terminology. We have used standard notation for number theory that can be found, for instance, in [7, 10]. Given a positive function $g(x)$, the big-O function of $g(x)$, denoted by $O(g(x))$, is a function such that $|O(g(x))|\leq Cg(x)$ uniformly. Observe that necessarily $C>0$, otherwise $O(g(x))$ vanishes, but $O(g(x))$ could be negative, positive or both. Moreover we can write $O(g(x))=C(x)g(x)$ for some uniformly bounded function $C(x)$. Throughout the paper we consider $x$ as a positive real variable, $p$ denotes prime numbers and $n$ and $d$ denote natural numbers. Prime zeta function is defined as the sum of the powers of reciprocals of prime numbers, i.e. $P(s)=\sum_{p}\frac{1}{p^{s}}$, where $s$ is a complex number, (see for instance [1, 4]), also $P_{x}(s)=\sum_{p\leq x}\frac{1}{p^{s}}$. The von Mangoldt function $\Lambda(n)$ is the arithmetic function $\Lambda(n)=\log p$ if $n=p^{k}$ for some positive integer $k$, $\Lambda(n)=0$ otherwise. The so-called Chebyshev functions are $\psi(x)=\sum_{n\leq x}\Lambda(n)$ and $\theta(x)=\sum_{p\leq x}\log p$, and the prime counting function is $\pi(x)=\sum_{p\leq x}1$. We define $N_{x}=\prod_{p\leq x}p$ and $\varphi_{x}=\prod_{p\leq x}(p-1)$. Observe that $\theta(x)=\log N_{x}$ and $\varphi_{x}=\varphi(N_{x})$, where $\varphi(n)=|\{m\in\mathbb{N}:m\leq n,(m,n)=1\}|$ is the Euler totient function. The floor function is $x]=\max\{m\in\mathbb{Z}:m\leq x\}$, so we can write $x]=x-\{x\}$, where function $\{x\}$ is bounded between 0 and $1$.
2. On the sign of the error term
In this unique section we prove a theorem and discuss the error term. The first one relates the sign of the error term of Mertens theorem for the sum of reciprocal of primes with the verification of the Nicolas inequality. To this aim we apply results involving the zeta prime function, basic estimates for Chevyshev functions and standard techniques of real analysis. In the second part we analyze the sign of such an error term, to this end we revise the proof of Mertens Theorem and obtain an explicit formula for the term, then we approximate sufficiently its value.
Theorem 2.1.
Suppose that $x\geq 2$, and let the Mertens formula
$\sum_{p\leq x}\frac{1}{p}=\log\log x+b+O(1/\log x)\,.$
(I) If there are $x_{0}\geq 2$ and $C>0$ such that $O(1/\log x)>\frac{C}{\log x}$ for every $x\geq x_{0}$, then (T) holds.
(II) If there are $x_{0}\geq 2$ and $C\leq 0$ such that $O(1/\log x)\leq\frac{C}{\log x}$ for every $x\geq x_{0}$, then (F) holds.
Proof.
We divide the proof in three steps.
Step 1. With the Taylor series of the logarithm we get
$\begin{split}\log\varphi_{x}&=\log\prod_{p\leq x}(p-1)=\log\Big(\prod_{p\leq x}p\prod_{p\leq x}\big(1-\frac{1}{p}\big)\Big)\\ &=\log\prod_{p\leq x}p+\log\prod_{p\leq x}\big(1-\frac{1}{p}\big)=\theta(x)-\sum_{p\leq x}\sum_{n=1}^{\infty}\frac{1}{np^{n}}\\ &=\theta(x)-\sum_{n=1}^{\infty}\frac{1}{n}\sum_{p\leq x}\frac{1}{p^{n}}=\theta(x)-\sum_{n=1}^{\infty}\frac{P_{x}(n)}{n}\,.\end{split}$
Taking into account Mertens Theorem (see e.g. [7, Theorem 8.8(d)]) and the above equation, we can rewrite the Nicolas inequality, namely $e^{\gamma}\log\log N_{x}<\dfrac{N_{x}}{\varphi_{x}}$, as
(2) $\begin{split}\gamma&<\log\Big(\frac{N_{x}}{\varphi_{x}\log\log N_{x}}\Big)=\log N_{x}-\log\varphi_{x}-\log\log\log N_{x}\\ &=\sum_{n=1}^{\infty}\frac{P_{x}(n)}{n}-\log\log\theta(x)=\sum_{n=2}^{\infty}\frac{P_{x}(n)}{n}-\log\log\theta(x)+\sum_{p\leq x}\frac{1}{p}\\ &=\sum_{n=2}^{\infty}\frac{P_{x}(n)}{n}-\log\log\theta(x)+\log\log x+b+O(1/\log x)\\ &=\sum_{n=2}^{\infty}\frac{P_{x}(n)}{n}+b+\log\Big(\frac{\log x}{\log\theta(x)}\Big)+O(1/\log x)\,.\end{split}$
Observe that $P_{x}(n)$ is increasing in $x$ and converges to $P(n)$, thus for each $\varepsilon>0$ there exists $x_{1}$ such that $P(n)-P_{x}(n)<\varepsilon$ for every $x\geq x_{1}$. In addition one can deduce that
(3) $\sum_{n=2}^{\infty}\dfrac{P(n)}{n}=\gamma-b\,$,
where $\gamma$ is the Euler constant and $b$ is known as the Mertens constant ([9, (2.7)]).
Step 2: (I). Let $\alpha>0$, as $C>0$ and $\log(1+\frac{\alpha}{\log x})$ is strictly decreasing and converges to zero, there exists $x_{2}$ such that $C>\log(1+\frac{\alpha}{\log x})$ for all $x\geq x_{2}$, hence by definition of the number $e$ as an increasing limit
$\begin{split}e^{C/\log x}&>1+\frac{C}{\log x}>1+\frac{\log(1+\frac{\alpha}{\log x})}{\log x}\\ &=\frac{\log x+\log(1+\frac{\alpha}{\log x})}{\log x}\\ &=\frac{\log(x(1+\frac{\alpha}{\log x}))}{\log x}=\log_{x}\Big(x\big(1+\frac{\alpha}{\log x}\big)\Big)\,.\end{split}$
Therefore $x^{e^{C/\log x}}>x(1+\frac{\alpha}{\log x})$ and for a suitable $\alpha$ we can ensure that $x^{e^{C/\log x}}>\theta(x)$ for every $x\geq x_{2}$, we can choose for instance $\alpha=1/2$ (see [9, Theorem 4(3.15)]). In consequence, taking logarithms twice at both sides, we first get $e^{C/\log x}\log x>\log\theta(x)$, and then $\log\Big(\frac{\log x}{\log\theta(x)}\Big)+\frac{C}{\log x}>0$ for every $x\geq x_{2}$. This implies that there exists some $\varepsilon>0$ such that
$\log\Big(\frac{\log x}{\log\theta(x)}\Big)+\frac{C}{\log x}-\varepsilon>0\,.$
Now we take into account computations of (2), equation (3), comments above it and add the Euler constant $\gamma$ to both sides of this inequality for $x\geq x_{3}=\max\{x_{0},x_{1},x_{2}\}$
$\begin{split}\gamma&<\gamma+\log\Big(\frac{\log x}{\log\theta(x)}\Big)+\frac{C}{\log x}-\varepsilon\\ &<\gamma-b+b+\log\Big(\frac{\log x}{\log\theta(x)}\Big)+O(1/\log x)-\varepsilon\\ &=\sum_{n=2}^{\infty}\frac{P(n)}{n}-\varepsilon-\log\log\theta(x)+\sum_{p\leq x}\frac{1}{p}\\ &<\sum_{n=2}^{\infty}\frac{P_{x}(n)}{n}-\log\log\theta(x)+\sum_{p\leq x}\frac{1}{p}=\log\Big(\frac{N_{x}}{\varphi_{x}\log\log N_{x}}\Big)\,.\end{split}$
Therefore Nicolas inequality holds for every $x\geq x_{3}$, which is not exactly (T), but it is in contradiction with (F), and as by [6, Theorem 2] there are only two possibilities, (T) must hold.
Step 3: (II). Now $C\leq 0$ and by hypothesis there is $x_{0}$ such that $O(1/\log x)<\frac{C}{\log x}\leq 0$ for every $x\geq x_{0}$, moreover we know there is some $x_{4}\geq x_{0}$ such that $\theta(x_{4})>x_{4}$ (see, for instance, the first page in [8]), join these facts with reasonings in previous steps, we obtain
[…]
hence Nicolas inequality does not hold for $x_{4}$ and we can apply the same argument to every integer $x\geq x_{0}$ such that $\theta(x)>x$ which is an infinite set of integers, i.e. assertion (F) holds. ∎
Observe that the failure of Nicolas inequality depends directly from the change of sign of $\theta(x)-x$. In the following subsection we have used the quick guide of Mertens Theorem given in [2].
2.1. Discussion on the sign of the error term
Step 1. Let start with an estimate for the sum of logarithms (see [7, Lemma 8.2 and Theorem 8.8(a)] and also [2]).
$\sum_{n\leq x}\log x=x\log x-x+O_{1}(\log x)\,$,
where do not mind the sign of $O_{1}(\log x)$ because in next step we will divide it by $x$.
Step 2. We follow with a sum involving the Mangoldt function (see [7, Theorem 8.8(a)]).
$\sum_{d\leq x}\frac{\Lambda(d)}{d}=\log x+O_{2}(1)\,$,
where
$O_{2}(1)=\frac{O_{1}(\log x)}{x}+\frac{\displaystyle\sum_{d\leq x}\Big\{\dfrac{x}{d}\Big\}\Lambda(d)}{x}\,.$
On the one hand we argue by computation that the second term of this sum is between $0.41$ and $0.43$ for sufficiently large $x$. This approximation has been computed with the formula
(4) $0.41<\frac{\displaystyle\sum_{n=1}^{800}\Big(\displaystyle\sum_{j=1}^{[\pi(k)^{1/n}]}\Big(\frac{k}{p_{j}^{n}}-\Big[\frac{k}{p_{j}^{n}}\Big]\Big)\log p_{j}\Big)}{k}<0.43\,$,
for $k\in(10^{4},10^{5})$ and a list of primes $\{p_{j}\}_{j}$. Alternatively we found estimates from above and from below for the second term. We can assume that $x$ is integer, hence the reminder of $\dfrac{x}{d}$ is $1$ for $d\in(x/2,x)$, so $\Big\{\dfrac{x}{d}\Big\}=\dfrac{1}{d}$ for $d\in(x/2,x)$, also recall that $\Big\{\dfrac{x}{x}\Big\}=0$. Taking into account [7, Theorem 8.3], we can assert that $\psi(x/2)<3/4\psi(x)$ for a sufficient large integer $x$, thus
$\begin{split}\sum_{d\leq x}\Big\{\frac{x}{d}\Big\}\Lambda(d)&=\sum_{d\leq x/2}\Big\{\frac{x}{d}\Big\}\Lambda(d)+\sum_{x/2<d<x}\frac{1}{d}\Lambda(d)\\ &<\sum_{d\leq x/2}\Lambda(d)+\frac{2}{x+1}\sum_{x/2<d<x}\Lambda(d)\\ &<\psi(x/2)+\frac{2}{x+1}(\psi(x)-\psi(x/2))\\ &<\frac{3}{4}\psi(x)+\frac{2}{x+1}\psi(x)=\frac{3x+11}{4x+4}\psi(x).\end{split}$
Moreover, if we avoid the terms for $d\leq x/2$ in the first equality, we have
$\sum_{d\leq x}\Big\{\frac{x}{d}\Big\}\Lambda(d)>\frac{1}{x}\big(\psi(x)-\psi(x/2)\big)\,.$
In consequence, taking into account that $ax<\psi(x)<cx$ for $a=0.780355$... and $c=1.170533$..., for a sufficiently large $x$ (see [7, Theorem 8.3]), we deduce that
$\frac{0.1950884}{x}<\frac{\displaystyle\sum_{d\leq x}\Big\{\dfrac{x}{d}\Big\}\Lambda(d)}{x}<0.58526625\,$,
where 0.1950884<ac/20.1950884<a-c/2 and $0.58526625<\dfrac{3}{4}a$. On the other hand, as $O_{1}(\log x)/x$ is negative and converges to zero, there is a $x_{2}$ such that $0<O_{2}(1)<0.585$ for every $x\geq x_{2}$.
Step 3. Next we deal with a sum over primes (see [7, Theorem 8.8(b)]).
$\sum_{p\leq x}\frac{\log p}{p}=\log x+O_{3}(1)\,$,
where
$O_{3}(1)=O_{2}(1)-\sum_{n=2}^{\infty}\sum_{p\leq x^{1/n}}\frac{\log p}{p^{n}}\,.$
We approximate the second term by means of a list of primes and the formula
(5) $\sum_{n=2}^{40}\left(\sum_{j=1}^{[\pi(10^{5})^{1/n}]}\frac{\log p_{j}}{p_{j}^{n}}\right)=0.753365132271.$..\,.
Let us obtain two coarse bounds. We know that the second term is bounded and, since $\log p/p^{n}$ is positive, it increases when the $x$ grows and the partial sums in $n$ grows. Moreover it is bounded from above by $\sum_{j=2}^{\infty}\frac{\log j}{j(j-1)}$ (see proof of [7, Theorem 8.8(b)]), which is bounded by the integral of $\dfrac{\log(x-1)}{(x-1)(x-2)}$ between $2$ and infinity. So, for sufficiently large $x$
$\begin{split}0.633&<\sum_{n=2}^{40}\Big(\sum_{j=1}^{[\pi(100)^{1/n}]}\frac{\log p_{j}}{p_{j}^{n}}\Big)<\sum_{n=2}^{\infty}\sum_{p\leq x^{1/n}}\frac{\log p}{p^{n}}<\sum_{j=2}^{\infty}\frac{\log j}{j(j-1)}\\ &<\int_{2}^{\infty}\frac{\log(j-1)}{(j-1)(j-2)}=\frac{\pi^{2}}{6}<1.645\,.\end{split}$
where $(p_{j})_{j}$ denotes the sequence of primes. This implies that there exists $x_{3}$ such that, for every $x\geq x_{3}$:
(6) $1.645<O_{3}(1)<-0.048<0.585-0.633$
Step 4. Finally we study the error term of the sum of reciprocal of primes (see [7, Theorem 8.8(d)]). We define $R(x)=\displaystyle\sum_{p\leq x}\dfrac{\log p}{p}-\log x$, so thanks to the estimate given in previous steps we get $R(x)=O_{3}(1)$. Let
$\sum_{p\leq x}\frac{1}{p}=\log\log x+b+O(1/\log x)\,$,
where
(7) $O(1/\log x)=\frac{R(x)}{\log x}-\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt\,.$
First observe that
$\frac{r}{\log x}-\int_{x}^{\infty}\frac{s}{t(\log t)^{2}}=\frac{r-s}{\log x}\,.$
Having in mind the above estimate for $O_{3}(1)=R(x)$, we assume that $r,s\in(-1.645,-0.048)$ and obtain that
$\frac{-1.597}{\log x}<O(1/\log x)<\frac{1.597}{\log x}\,$,
for every $x\geq x_{0}=\max\{x_{1},x_{2},x_{3}\}$.
Remark $2.$
Observe that from approximations (4) and (5) we have that
$|O(1/\log x)|<0.02/\log x\,$,
for sufficiently large $x$.
Remark 2.3.
Estimates above are trivial, but we have seen that $R(x)$ is decreasing almost everywhere (a.e.) and negative for sufficiently large $x$. Let the expression of $b$ given in proof of [3, Theorem 427] (see also [2, Propostion]), that is
$\begin{split}b&=\int_{2}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt+1-\log\log 2\\ &=\int_{2}^{x}\frac{R(t)}{t(\log t)^{2}}\,dt+1-\log\log 2+\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt\,.\end{split}$
Applying Stieltjes integration by parts a.e., we obtain
$\int_{2}^{x}\frac{R(t)}{t(\log t)^{2}}\,dt=\frac{R(2)}{\log 2}-\frac{R(x)}{\log x}+\int_{2}^{x}\frac{R^{\prime}(t)}{\log t}\,dt=-1-\frac{R(x)}{\log x}+\sum_{p\leq x}\frac{1}{p}+\int_{2}^{x}\frac{R^{\prime}(t)}{\log t}\,dt\,$,
Moreover $\int_{2}^{x}\frac{R(t)}{t(\log t)^{2}}\,dt$ is negative and strictly decreasing that converges to $b+\log\log 2-1$, hence
$\int_{2}^{x}\frac{R(t)}{t(\log t)^{2}}\,dt>b+\log\log 2-1\,$,
for sufficiently large $x$. Notice that $R(x)$ is derivable and integrable (a.e.) and so is its derivative, moreover $R(2)/\log 2=-1/2$. Therefore
$\begin{split}-\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt&=-b+1-\log\log 2+\int_{2}^{x}\frac{R(t)}{t(\log t)^{2}}\,dt\\ &=-b+1-\log\log 2-1-\frac{R(x)}{\log x}+\sum_{p\leq x}\frac{1}{p}+\int_{2}^{x}\frac{R^{\prime}(t)}{\log t}\,dt\,.\end{split}$
With these things in hand we see that
$\frac{R(x)}{\log x}-\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt=-b-\log\log 2+\sum_{p\leq x}\frac{1}{p}+\int_{2}^{x}\frac{R^{\prime}(t)}{\log t}\,dt\,.$
However, from these equations we just can deduce that $\int_{2}^{x}\frac{R^{\prime}(t)}{\log t}=\log\log 2-\log\log x$.
Remark 2.4.
Assume that
$O(1/\log x)=\frac{R(x)}{\log x}-\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt=\frac{C(x)}{\log x}\,$,
thus
$\begin{split}C(x)&=\log x\cdot O(1/\log x)=R(x)-\log x\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt\\ &=\sum_{p\leq x}\frac{\log p}{p}-\log x\Big(1+\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt\Big)\end{split}$
and $\int_{x}^{\infty}\frac{R(t)}{t(\log t)^{2}}\,dt$ goes to zero when $x$ grows to infinity. In consequence $C(x)$ goes to $R(x)$ that is negative and bounded from above by a negative number (see (6)), that is $C(x)<C$ for some $C<0$